Calculate the pearson correlation coefficient for the following data

X 7 2 8 4 5

Y 4 4 7 6 7


1
Expert's answer
2021-02-08T17:27:23-0500

The Pearson correlation coefficient formula is given below,

"r=\frac{n(\Sigma XY)-(\Sigma X)(\Sigma Y)}{\sqrt{[n\Sigma X^2-(\Sigma X)^2][n\Sigma Y^2-(\Sigma Y)^2]}}", where "n=" sample size

To find Pearson correlation coefficient of the given data, we first calculate the following table



Here "n=5" .

"\therefore" "r=\frac{n(\Sigma XY)-(\Sigma X)(\Sigma Y)}{\sqrt{[n\Sigma X^2-(\Sigma X)^2][n\Sigma Y^2-(\Sigma Y)^2]}}" "=\frac{(5×151)-(26×28)}{\sqrt{[(5×158)-(26)^2].[(5×166)-(28)^2]}}" "=\frac {27}{\sqrt {5244}}=0.37" (approximately)